OpenStudy (anonymous):

The radius of the cone is 1.25 inches, and its height is 2.75 inches. If the diameter of the bubble gum ball is 0.5 inches, what is the closest approximation of the volume of the cone that can be filled with flavored ice?

OpenStudy (lgbasallote):

uhhhh what is the relation of the bubble gum ball if it's just asking volume of cone -__- is this all the question??

OpenStudy (lgbasallote):

volume of cone = h/3 (pi r^2) V = (2.75/3)pi(1.25)^2 calculator would be nice here...

OpenStudy (anonymous):

Thats why I am confused also lol because this is the whole question....

OpenStudy (lgbasallote):

i think that bubble gum ball is a misleader..it does not relate to anything...

OpenStudy (anonymous):

Would the answer be 4.50in3

OpenStudy (lgbasallote):

pretty much